OpenStudy (anonymous):

Given an exponential function for compounding interest, A(x) = P(1.02)x, what is the rate of change?

OpenStudy (anonymous):

2% 0.02% 1.02% 102%

OpenStudy (jdoe0001):

have you covered the compound interest formula yet?

OpenStudy (anonymous):

No

OpenStudy (jdoe0001):

do you know how to change percentage quantities to decimal format?

OpenStudy (anonymous):

No im very lost on this one.

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%=\cfrac{{\color{red}{ a}}}{100} \qquad \qquad {\color{red}{ 7}}\%=\cfrac{{\color{red}{ 7}}}{100}\implies \textit{what do you think?}\)

OpenStudy (anonymous):

is it .02?

OpenStudy (jdoe0001):

7/100? is not 0.02

OpenStudy (anonymous):

I know its .07 i am talking about the answer

OpenStudy (jdoe0001):

ohh .. well.. how did you 0.02?

OpenStudy (jdoe0001):

\(\bf A(x)=P(1.02)^x\qquad \textit{compound formula}\implies A=P\left(1+\frac{r}{n}\right)^{nx} \\ \quad \\ n=periods\qquad x=years\qquad r=rate\textit{ in decimal format} \\ \quad \\ \textit{for 1 year and annual compounding, so 1 period per year} \\ \quad \\ A=P\left(1+\frac{r}{n}\right)^{nx}\implies A=P\left(1+\frac{r}{1}\right)^{1x}\implies A=P(1+r)^x\) so.. what do you think?

OpenStudy (anonymous):

This doesn't make sense. Its okay Thanks for trying!

OpenStudy (anonymous):

It's 2% you divide compound 100 then add to 1